OpenStudy (anonymous):

x=3y 3x-5y=12 solve, please?

OpenStudy (anonymous):

multiply top eqn by 3 gives 3x=9y

OpenStudy (anonymous):

so if 3x=9y we can put in bottom

OpenStudy (anonymous):

9y-5y=12

OpenStudy (anonymous):

12?

jimthompson5910 (jim_thompson5910):

3x-5y=12 3(3y)-5y=12 9y-5y=12 4y = 12 4y/4 =12/4 y = 3 --------------------- x = 3y x= 3(3) x =9 So the solutions are x = 9 and y = 3